# Option Pricing with Time-Varying Volatility Risk Aversion

### v3

@@ A.5 Proof of Theorem 5
such that
A.5.1 The Distribution of the Score
Define
Then the MGF of score $s_{t}$ can be expressed as
$\mathbb{E}_{t-1}^{\mathbb{P}}\left[\exp\left(us_{t}\right)\right]$ $=\mathbb{E}_{t-1}^{\mathbb{P}}\left[\exp\left(u\frac{\psi_{t}^{\prime}e_{t}}{\sqrt{\psi_{t}^{\prime}\Sigma_{N_{t}}\psi_{t}}}\right)\right]=\mathbb{E}_{t-1}^{\mathbb{P}}\left[\mathbb{E}_{t-1}^{\mathbb{P}}\left[\left.\exp\left(u\frac{\psi_{t}^{\prime}e_{t}}{\sqrt{\psi_{t}^{\prime}\Sigma_{N_{t}}\psi_{t}}}\right)\right|\psi_{t}\right]\right].$
Under Assumption 4 where $e_{t}\sim iid\ N(0,\Sigma_{N_{t}})$ the MGF simplifies to
$\mathbb{E}_{t-1}^{\mathbb{P}}\left[\left.\exp\left(u\frac{\psi_{t}^{\prime}e_{t}}{\sqrt{\psi_{t}^{\prime}\Sigma_{N_{t}}\psi_{t}}}\right)\right|\psi_{t}\right]$ $=\exp\left(\tfrac{1}{2}\frac{u\psi_{t}^{\prime}}{\sqrt{\psi_{t}^{\prime}\Sigma_{N_{t}}\psi_{t}}}\Sigma_{N_{t}}\frac{u\psi_{t}}{\sqrt{\psi_{t}^{\prime}\Sigma_{N_{t}}\psi_{t}}}\right)=\exp\left(\tfrac{1}{2}u^{2}\right).$
Therefore, the score, $s_{t}$, is iid with a standard normal distribution under $\mathbb{P}$.
To derive the distribution under $\mathbb{Q}$, we first obtain,
$\mathbb{E}_{t}^{\mathbb{P}}\left[\exp\left(us_{t+1}+\phi_{t}R_{t+1}+\xi_{t}h_{t+2}\right)\right]$
$=$ $\mathbb{E}_{t}^{\mathbb{P}}\left\{\mathbb{E}_{t}^{\mathbb{P}}\left[\exp\left(us_{t+1}+\phi_{t}R_{t+1}+\xi_{t}h_{t+2}\right)|\phi_{t}R_{t+1},\xi_{t}h_{t+2},\psi_{t+1}\right]\right\}$
$=$ $\mathbb{E}_{t}^{\mathbb{P}}\left\{\mathbb{E}_{t}^{\mathbb{P}}\left[\exp\left(\tfrac{1}{2}u^{2}+\phi_{t}R_{t+1}+\xi_{t}h_{t+2}\right)|\phi_{t}R_{t+1},\xi_{t}h_{t+2},\psi_{t+1}\right]\right\}$
where the last equality is a consequence of $e_{t}$ being assumed to be independent of other variables. Next, we have
$\mathbb{E}_{t}^{\mathbb{Q}}\left(\exp\left(us_{t+1}\right)\right)$ $=\frac{\mathbb{E}_{t}^{\mathbb{P}}\left[\exp\left(us_{t+1}+\phi_{t}R_{t+1}+\xi_{t}h_{t+2}\right)\right]}{\mathbb{E}_{t}^{\mathbb{P}}[\exp(\phi_{t}R_{t+1}+\xi_{t}h_{t+2})]}=\exp\left(\tfrac{1}{2}u^{2}\right),$
A.6 Proof of Theorem 3
Let $\boldsymbol{\eta}_{t,M}=\left(\log\eta_{t+1},\ldots,\log\eta_{t+M}\right)^{\prime}$ and $\bar{\boldsymbol{\eta}}_{t,M}^{e}=\mathbb{E}_{t}^{\mathbb{Q}}(\boldsymbol{\eta}_{t,M})$, where the latter embeds the expected path of $\log\eta_{t}$. The MGF for $M$-period cumulative returns $R_{t,M}$ can be expressed as
@@ A.6 Proof of Theorem 3
where $f(\boldsymbol{\eta}_{t,M})\equiv\mathbb{E}_{t}^{\mathbb{Q}}[\exp(sR_{t,T})\mid\boldsymbol{\eta}_{t,M}]$. A second-order Taylor expansion of $f$ around $\bar{\boldsymbol{\eta}}_{t,M}^{e}$ gives us,
@@
is the Hessian, $\nabla$ is the Jacobian, and $\bm{\varepsilon}_{t,M}=\boldsymbol{\eta}_{t,M}-\bar{\boldsymbol{\eta}}_{t,M}^{e}$. Since $\bar{\boldsymbol{\eta}}_{t,M}^{e}\in\mathcal{G}_{t}$ and $f(\bar{\boldsymbol{\eta}}_{t,M}^{e})=g_{t,M}(s)$ we have
$g_{t,M}(s)\approx g_{t,M}(s|\bar{\boldsymbol{\eta}}_{t,M}^{e})+\frac{1}{2}\mathrm{tr}\left\{\tilde{H}_{t,M}(s)\Sigma_{M}\right\},$ (A.11)
@@
A.6.1 The Hessian Matrix $\tilde{H}_{t,M}(s)$
@@
$A(M)\equiv A_{T}(s,M),$ and $B(M)\equiv B_{T}(s,M)$. For derivatives of a variable $Y$ we write
@@
From Theorem 2 we have $f(\boldsymbol{\eta}_{t,M})=\exp\left[A(M)+B(M)h_{t+1}^{*}\right]$ with first and second derivatives given by
@@
$\omega_{m}^{(i)}=\begin{cases}\omega_{m}&\text{for }i=m,\\ 0&\text{otherwise},\end{cases}\quad\alpha_{m}^{(i)}=\begin{cases}\alpha_{m}&\text{for }i=m,m+1\\ 0&\text{otherwise,}\end{cases},$
$\beta_{m}^{(i)}=\begin{cases}\text{$$}\beta_{m}&\text{for }i=m\\ -\beta_{m}&\text{for }i=m+1\\ 0&\text{otherwise,}\end{cases},\quad\gamma_{m}^{(i)}=\begin{cases}\gamma_{m}-\tfrac{1}{2}&\text{for }i=m+1\\ 0&\text{otherwise,}\end{cases},$
$\omega_{m}^{(i,j)}=\begin{cases}\omega_{m}&\text{for }i=j=m\\ 0&\text{otherwise,}\end{cases},\quad\alpha_{m}^{(i,j)}=\begin{cases}\alpha_{m}&\text{for }i,j=m,m+1\\ 0&\text{otherwise,}\end{cases},$
$\beta_{m}^{(i,j)}=\begin{cases}\text{$$}\beta_{m}&\text{for }i=j\in\{m,m+1\}\\ -\beta_{m}&\text{for }i\neq j\in\{m,m+1\}\\ 0&\text{otherwise,}\end{cases},\quad\gamma_{m}^{(i,j)}=\begin{cases}\gamma_{m}-\tfrac{1}{2}&\text{for }i=j=m+1\\ 0&\text{otherwise.}\end{cases}$
@@
A.7 The Scores for Derivatives
A.7.1 Score for VIX
@@
$\frac{\partial a_{1}(M,\sigma^{2})}{\partial\log\eta_{t}}$ $=\frac{A^{2}}{M}\sum_{k=2}^{M}\sum_{i=2}^{k}\tilde{\beta}^{k-i}\left[\omega\mathbb{E}_{t}^{\mathbb{Q}}(\eta_{t+k-1})B_{k-1,1}+\alpha\mathbb{E}_{t}\left(\eta_{t+k-1}\eta_{t+i-2}\right)\left(B_{k-1,1}+B_{i-2,1}\right)\right]$
@@
A.7.2 Score for Option Prices
When the model is estimated with option prices, the elements of $X_{t}\in\mathbb{R}^{N_{t}\times 1}$ are given by
$\log\left({\rm IV}_{bs}\left(C_{t},S_{t},M,K,r\right)\right),$
for the $N_{t}$ options selected at time $t$. The logarithm of Black-Scholes implied volatility makes this variable comparable to $\log\mathrm{VIX}_{t}$, which is used when the model is estimated with VIX data.
@@ A.7.2 Score for Option Prices
A.8 Computing Moments by the Moment-Generating function
The $k$-th moment of cumulative returns can be expressed as
We have here used the following Laplace transformations for $R^{k}$:
$\int_{-\infty}^{0}R^{k}e^{-uR}\mathrm{d}R$ $=$ $-\frac{k!}{u^{k+1}},\quad u_{R}<0,$
Thus, it follows that
@@ A.9 Computation of Option Prices by Numerical Integrations
Appendix B Supplementary Empirical Results
B.1 Alternative RMSE Measure
@@
${\rm RMSE_{VIX}}=\sqrt{\frac{1}{T}\sum_{t=1}^{T}\text{$\left[\mathrm{VIX}_{t}^{m}-\mathrm{VIX}_{t}\right]^{2}$}},$ (B.1)
@@
${\rm RMSE_{IV}}=\sqrt{\frac{1}{\sum_{t=1}^{T}N_{t}}\sum_{t=1}^{T}\text{$\sum_{i=1}^{N_{t}}\left[\mathrm{IV}_{t,i}^{m}-\mathrm{IV}_{t,i}\right]^{2}$}}\times 100.$ (B.2)
The results for the RMSE in volatility levels are reported in Tables B.1 and B.2, where the former is analogous to Table 4 and the latter is analogues to the out-of-sample results in Table 5.
Table B.1: VIX and Option Pricing Performance (RMSE)
@@
Table B.2: Out-of-sample VIX and Option Pricing (RMSE)
@@
The results are qualitatively very similar to the RMSE for logarithmically transformed volatilities in Tables 4 and 5.
@@
B.2 A Variance Risk Ratio based on Model-Free Realized Variances
@@
As a robustness check, we compute an alternative monthly measure of the VRR, $\tilde{\eta}=\frac{\mathrm{VIX}^{2}}{\sigma_{\mathbb{P}}^{2}}$, where $\sigma_{\mathbb{P}}^{2}$ is an empirical measure of the expected variance under $\mathbb{P}$, deduced from monthly realized variances. So, $\sigma_{\mathbb{P}}^{2}$ does not rely on the Heston-Nandi GARCH model. The monthly realized variance for the S&P 500 index is computed as the sum of squared 5-minute intraday returns during the hours from 9:30 AM to 4:00 PM plus the sum of squared overnight returns. This quantity is multiplied by $12\times 100^{2}$ to bring it to the same scale as the $\mathrm{VIX}^{2}$ (squared annualized percentage). Finally, $\sigma_{\mathbb{P}}^{2}$ is defined to be the one-month ahead prediction of the monthly realized variance using a AR(1) model.
@@
Figure B.1: An empirical monthly variance risk ratio, defined as the ratio of ${\rm VIX}_{t}^{2}$ and the predicted value of the monthly realized variance using an AR(1) model.
@@
This time-series of this empirical VRR is presented in Figure B.1, and it is very similar to the model-based time-series presented in Figure 5.
@@
Figure B.2 contains scatterplots of volatility prediction errors under $\mathbb{P}$ and the level of $\log\eta$. The former is defined as the difference between the ex-post model-implied variance and the ex-ante model-based one-month ahead volatility prediction. The scatterplots indicate some positive correlation, but we do not find that a large proportion of overpredictions of volatility occurs when $\log\eta_{t}$ is relatively small, nor that a large proportion of underpredictions of volatility occurs when $\log\eta_{t}$ is large.
@@
Figure B.2: The relations between DHNG model-implied logarithm of variance risk ratio and one-month ahead volatility prediction errors.
B.3 Empirical Results for Auxiliary Structure (Predetermined VRR)
@@
While the Auxiliary structure does not constitute a coherent option pricing model, we can estimate the parameters. We should not expect all parameters to be consistent, because the auxiliary structure employs and option pricing that poorly approximates the true formula, as shown in Table 1. The parameter estimates of the Auxiliary structure are shown in Table B.3, where we also include those of the DHNG model, for comparison. Most parameter estimates are very similar, with the important exception with the intercept in the score-driven model for $\log\eta_{t}$. This can be seen from the estimate of $\zeta=\mathbb{E}\left(\log\eta_{t}\right)$, which is substantially larger for the Auxiliary structure than the DHNG model. The reason is simple that the Auxiliary structure need an upwards biased value of $\eta_{t}$ to compensate for the shortcomings of its option pricing formula, that neglects the random variations in $\eta_{t}$. So, by inflating the model-implied value of $\eta_{t}$, the Auxiliary structure is able to reduce option pricing errors, which is part of the objective in the estimation problem.
@@
The option pricing performance for the Auxiliary structure is shown in Table B.4, where we, for the sake of comparison, include the results for the DHNG model. We evaluate the derivative pricing in terms of the RMSE for both the logarithmically transformed implied volatility (the first two columns) and the level of implied volatility (the last two columns). Interestingly, the Auxiliary structure only results in slightly larger pricing errors than the DHNG model. This can be ascribed to the adaptive nature of the score-driven model, which shifts $\log\eta_{t}$ to a higher level to compensate for errors embedded in the Auxiliary pricing formulae. Thus, while the first-order approximation is insufficient, the estimated score-drive model “rescues” the Auxiliary structure, by inflating the variance risk ratio above its true value.
Table B.3: Estimation Results for Auxiliary and DHNG
@@
Table B.4: Pricing Performance for Auxiliary and DHNG
@@
B.4 Supplementary Material for Table 6
@@
Economic Policy Uncertainty 0.258*** 0.203***
@@
Survey-based Uncertainty 0.106*** 0.065***
@@
Economic Uncertainty Index 0.037*** -0.007
@@
Variance Risk Premium 0.097*** 0.074***
@@
Stock Market Volatility 0.317*** 0.318*** 0.299*** 0.300*** 0.239*** 0.297*** 0.290*** 0.202*** 0.148***
@@
Default Yield Spread 0.162*** 0.108** 0.132*** 0.143*** 0.198*** 0.093* 0.080 0.160*** 0.100**
@@
Default Return Spread 0.031*** 0.031*** 0.029*** 0.028*** 0.026*** 0.029*** 0.027*** 0.024*** 0.021***
@@
Long Term Yield 0.040*** 0.046*** 0.027*** 0.027*** 0.056*** 0.021** 0.035*** 0.033*** 0.033***
@@
Dividend Price Ratio -0.158** -0.158** -0.118* -0.135* -0.233*** -0.116* -0.165** -0.174*** -0.179***
@@
Earnings Price Ratio -0.052 -0.031 -0.033 -0.019 -0.069* -0.073* -0.001 -0.004 -0.014
@@
Book to Market Ratio 0.014 -0.318 -0.184 -0.136 -0.158 0.040 0.071 -0.042 -0.566***
@@
Long Term Rate of Return 0.007* 0.007** 0.006 0.006 0.005 0.005 0.005 0.005 0.003

### v4

@@ A.5 Proof of Theorem 3
Let $\boldsymbol{\eta}_{t,M}=\left(\log\eta_{t+1},\ldots,\log\eta_{t+M}\right)^{\prime}$ and $\bar{\boldsymbol{\eta}}_{t,M}^{e}=\mathbb{E}_{t}^{\mathbb{Q}}(\boldsymbol{\eta}_{t,M})$ which embeds the expected path of $\log\eta_{t}$. The MGF for $M$-period cumulative returns $R_{t,M}$ can be expressed as
@@
where $f(\boldsymbol{\eta}_{t,M})\equiv\mathbb{E}_{t}^{\mathbb{Q}}[\exp(sR_{t,T})|\boldsymbol{\eta}_{t,M}]$. A second-order Taylor expansion of $f(\boldsymbol{\eta}_{t,M})$ around $\bar{\boldsymbol{\eta}}_{t,M}^{e}$ gives us,
@@
is the Hessian, $\nabla$ is the Jacobian, and $\bm{\varepsilon}_{t,M}=\boldsymbol{\eta}_{t,M}-\bar{\boldsymbol{\eta}}_{t,M}^{e}$. Since $\bar{\boldsymbol{\eta}}_{t,M}^{e}\in\mathcal{G}_{t}$ and $f(\bar{\boldsymbol{\eta}}_{t,M}^{e})=g_{t,M}(s|\bar{\boldsymbol{\eta}}_{t,M}^{e})$, by taking conditional expectations on both sides, we arrive
$g_{t,M}(s)\approx g_{t,M}(s|\bar{\boldsymbol{\eta}}_{t,M}^{e})+\frac{1}{2}\mathrm{tr}\left\{\tilde{H}_{t,M}(s)\Sigma_{M}\right\},$
@@
A.5.1 The Hessian Matrix $\tilde{H}_{t,M}(s)$
@@
and $A(M)\equiv A_{T}(s,M),$ $B(M)\equiv B_{T}(s,M)$. For derivatives of a variable $Y$ we write
@@
From Theorem 2, we have $f(\boldsymbol{\eta}_{t,M})=\exp\left[A(M)+B(M)h_{t+1}^{*}\right]$ with first and second derivatives given by
@@
$\omega_{m}^{(i)}=\begin{cases}\omega_{m}&\text{for }i=M-m,\\ 0&\text{otherwise},\end{cases}\quad\alpha_{m}^{(i)}=\begin{cases}\alpha_{m}&\text{for }i=M-m,M-m-1\\ 0&\text{otherwise,}\end{cases},$
$\beta_{m}^{(i)}=\begin{cases}\text{$$}\beta_{m}&\text{for }i=M-m\\ -\beta_{m}&\text{for }i=M-m-1\\ 0&\text{otherwise,}\end{cases},\quad\gamma_{m}^{(i)}=\begin{cases}\tfrac{1}{2}-\gamma_{m}&\text{for }i=M-m-1\\ 0&\text{otherwise,}\end{cases},$
$\omega_{m}^{(i,j)}=\begin{cases}\omega_{m}&\text{for }i=j=M-m\\ 0&\text{otherwise,}\end{cases},\quad\alpha_{m}^{(i,j)}=\begin{cases}\alpha_{m}&\text{for }i,j=M-m,M-m-1\\ 0&\text{otherwise,}\end{cases},$
$\beta_{m}^{(i,j)}=\begin{cases}\text{$$}\beta_{m}&\text{for }i=j\in\{M-m,M-m-1\}\\ -\beta_{m}&\text{for }i\neq j\in\{M-m,M-m-1\}\\ 0&\text{otherwise,}\end{cases},\quad\gamma_{m}^{(i,j)}=\begin{cases}\gamma_{m}-\tfrac{1}{2}&\text{for }i=j=M-m-1\\ 0&\text{otherwise.}\end{cases}$
@@
A.6 Proof of Theorem 4
Next we prove Theorem 4 that enables us to computing moments from the moment-generating function The $k$-th moment of cumulative returns can be expressed as
We have here used the following Laplace transformations for $R^{k}$:
$\int_{-\infty}^{0}R^{k}e^{-uR}\mathrm{d}R$ $=$ $-\frac{k!}{u^{k+1}},\quad u_{R}<0,$
Thus, it follows that
More general results for moment deduced by integrating the moment-generating function can be found in Hansen and Tong, (2024).
A.7 Proof of Theorem 5
such that
A.7.1 The Distribution of the Score
Define
Recall that $\psi_{t}^{\prime}$ is $\mathcal{G}_{t}$measurable, then under Assumption 4 where $e_{t}|\mathcal{G}_{t}\sim iid\ N(0,\Sigma_{N_{t}})$, the conditional MGF of score $s_{t}$ can be expressed as
$\mathbb{E}_{t}^{\mathbb{P}}\left[\exp\left(us_{t}\right)\right]$ $=\mathbb{E}_{t}^{\mathbb{P}}\left[\exp\left(u\frac{\psi_{t}^{\prime}e_{t}}{\sqrt{\psi_{t}^{\prime}\Sigma_{N_{t}}\psi_{t}}}\right)\right]=\exp\left(\tfrac{1}{2}\frac{u\psi_{t}^{\prime}}{\sqrt{\psi_{t}^{\prime}\Sigma_{N_{t}}\psi_{t}}}\Sigma_{N_{t}}\frac{u\psi_{t}}{\sqrt{\psi_{t}^{\prime}\Sigma_{N_{t}}\psi_{t}}}\right)=\exp\left(\tfrac{1}{2}u^{2}\right).$
Therefore, the scaled score, $s_{t}$, is iid with a standard normal distribution under $\mathbb{P}$.
To derive the distribution under $\mathbb{Q}$, we first obtain,
$\mathbb{E}_{t}^{\mathbb{P}}\left[\exp\left(us_{t}+\phi_{t}R_{t+1}+\xi_{t}h_{t+2}\right)\right]$
$=$ $\mathbb{E}_{t}^{\mathbb{P}}\left\{\mathbb{E}_{t}^{\mathbb{P}}\left[\exp\left(us_{t}+\phi_{t}R_{t+1}+\xi_{t}h_{t+2}\right)|\phi_{t}R_{t+1},\xi_{t}h_{t+2}\right]\right\}$
$=$ $\mathbb{E}_{t}^{\mathbb{P}}\left\{\mathbb{E}_{t}^{\mathbb{P}}\left[\exp\left(\tfrac{1}{2}u^{2}+\phi_{t}R_{t+1}+\xi_{t}h_{t+2}\right)|\phi_{t}R_{t+1},\xi_{t}h_{t+2}\right]\right\}$
where the last equality is a consequence of $e_{t}$ being assumed to be independent of any return shock. Then, we directly have
$\mathbb{E}_{t}^{\mathbb{Q}}\left(\exp\left(us_{t}\right)\right)$ $=\frac{\mathbb{E}_{t}^{\mathbb{P}}\left[\exp\left(us_{t}+\phi_{t}R_{t+1}+\xi_{t}h_{t+2}\right)\right]}{\mathbb{E}_{t}^{\mathbb{P}}[\exp(\phi_{t}R_{t+1}+\xi_{t}h_{t+2})]}=\exp\left(\tfrac{1}{2}u^{2}\right),$
A.8 The Scores for Derivatives
A.8.1 Score for VIX
@@
$\frac{\partial a_{1}(M,\sigma^{2})}{\partial\log\eta_{t}}$ $=\frac{A^{2}}{M}\sum_{k=2}^{M}\sum_{i=2}^{k}\tilde{\beta}^{k-i}\left[\omega\mathbb{E}_{t}^{\mathbb{Q}}(\eta_{t+k-1})B_{k-1,1}+\alpha\mathbb{E}_{t}^{\mathbb{Q}}\left(\eta_{t+k-1}\eta_{t+i-2}\right)\left(B_{k-1,1}+B_{i-2,1}\right)\right]$
@@
A.8.2 Score for Option Prices
When the model is estimated with option prices, the elements of $X_{t}\in\mathbb{R}^{N_{t}\times 1}$ are given by $\log\left({\rm IV}_{bs}\left(C_{t},S_{t},M,K,r\right)\right)$, for the $N_{t}$ options selected at time $t$. The logarithm of Black-Scholes implied volatility makes this variable comparable to $\log\mathrm{VIX}_{t}$, which is used when the model is estimated with VIX data.
@@ A.8.2 Score for Option Prices
When the number of daily available option prices, $N_{t}$, is large, computing the derivatives $\partial\log\text{IV}_{t,i}^{m}/\partial\log\eta_{t}$ for $i=1,2,\ldots,N_{t}$ becomes very time-consuming. In this case, one could assume that $\partial\log\text{IV}_{t,i}^{m}/\partial\log\eta_{t}\approx\chi_{t}>0$ (which is only $t$-dependent), such that the scaled score $s_{t}$, defined by (9) and Theorem 5, is approximated by
$s_{t}^{\text{appro}}=\frac{1}{\sigma_{e}}\frac{\iota_{N_{t}}^{\prime}\Omega_{N_{t}}^{-1}e_{t}}{\sqrt{\iota_{N_{t}}^{\prime}\Omega_{N_{t}}^{-1}\iota_{N_{t}}}}$
where $\iota_{N_{t}}$ is a $N_{t}\times 1$ vector of ones. With this approximation, there is no need to compute any derivatives. In our empirical analysis, we find the correlation between $s_{t}^{\text{appro}}$ and the true $s_{t}$ reaches up to 99.11%, such that the correlation between $\eta_{t}^{\text{appro}}$ and the true $\eta_{t}$ is 99.98%.
@@ A.9 Computation of Option Prices by Numerical Integrations
Online Appendix
S.1 Alternative RMSE Measure
@@
${\rm RMSE_{VIX}}=\sqrt{\frac{1}{T}\sum_{t=1}^{T}\text{$\left[\mathrm{VIX}_{t}^{m}-\mathrm{VIX}_{t}\right]^{2}$}},$ (S.1)
@@
${\rm RMSE_{IV}}=\sqrt{\frac{1}{\sum_{t=1}^{T}N_{t}}\sum_{t=1}^{T}\text{$\sum_{i=1}^{N_{t}}\left[\mathrm{IV}_{t,i}^{m}-\mathrm{IV}_{t,i}\right]^{2}$}}\times 100.$ (S.2)
The results for the RMSE in volatility levels are reported in Tables S.1 and S.2, where the former is analogous to Table 4 and the latter is analogues to the out-of-sample results in Table 5. The results are qualitatively very similar to the RMSE for logarithmically transformed volatilities in Tables 4 and 5.
Table S.1: VIX and Option Pricing Performance (RMSE)
@@
Table S.2: Out-of-sample VIX and Option Pricing (RMSE)
@@
S.2 A Variance Risk Ratio based on Model-Free Realized Variances
@@
As a robustness check, we compute an alternative monthly measure of the VRR, $\tilde{\eta}=\frac{\mathrm{VIX}^{2}}{\sigma_{\mathbb{P}}^{2}}$, where $\sigma_{\mathbb{P}}^{2}$ is an empirical measure of the expected variance under $\mathbb{P}$, deduced from monthly realized variances. So, $\sigma_{\mathbb{P}}^{2}$ does not rely on the Heston-Nandi GARCH model. The monthly realized variance for the S&P 500 index is computed as the sum of the squared daily close-to-close log returns within the same month. This quantity is multiplied by $12\times 100^{2}$ to bring it to the same scale as the $\mathrm{VIX}^{2}$ (squared annualized percentage). Finally, $\sigma_{\mathbb{P}}^{2}$ is defined to be the one-month ahead prediction of the monthly realized variance using a AR(1) model.
@@
Figure S.1: An empirical monthly variance risk ratio, defined as the ratio of ${\rm VIX}_{t}^{2}$ and the predicted value of the monthly realized variance using an AR(1) model.
Alt text: Time series of the empirical monthly variance risk ratio, calculated as the ratio of VIX squared to the predicted monthly realized variance, where the prediction is based on a simple first-order autoregressive model.
@@
This time-series of this empirical VRR is presented in Figure S.1, and it is very similar to the model-based time-series presented in Figure 5.
@@
Figure S.2 contains scatterplots of one-month-ahead volatility prediction errors under $\mathbb{P}$ and the level of $\log\eta$. The former is defined as the difference between the ex-post model-implied variance and the ex-ante model-based one-month ahead volatility prediction. The scatterplots indicate some positive correlation, but we do not find that a large proportion of overpredictions of volatility occurs when $\log\eta_{t}$ is relatively small, nor that a large proportion of underpredictions of volatility occurs when $\log\eta_{t}$ is large.
@@
Figure S.2: Volatility prediction errors for one-month ahead predictions plotted against the variance risk ratio. Based on the DHNG model estimated with option prices (left panel) and VIX (right panel).
Alt text: Volatility prediction errors for one-month-ahead forecasts plotted against the variance risk ratio. Predictions are based on the DHNG model estimated using either option prices or VIX data.
S.3 Empirical Results for Auxiliary Structure (Predetermined VRR)
@@
While the Auxiliary structure does not constitute a coherent option pricing model, we can estimate the parameters. We should not expect all parameters to be consistent, because the auxiliary structure employs and option pricing that poorly approximates the true formula, as shown in Table 1. The parameter estimates of the Auxiliary structure are shown in Table S.3, where we also include those of the DHNG model, for comparison. Most parameter estimates are very similar, with the important exception with the intercept in the score-driven model for $\log\eta_{t}$. This can be seen from the estimate of $\zeta=\mathbb{E}\left(\log\eta_{t}\right)$, which is substantially larger for the Auxiliary structure than the DHNG model. The reason is simple that the Auxiliary structure need an upwards biased value of $\eta_{t}$ to compensate for the shortcomings of its option pricing formula, that neglects the random variations in $\eta_{t}$. So, by inflating the model-implied value of $\eta_{t}$, the Auxiliary structure is able to reduce option pricing errors, which is part of the objective in the estimation problem.
@@
The option pricing performance for the Auxiliary structure is shown in Table S.4, where we, for the sake of comparison, include the results for the DHNG model. We evaluate the derivative pricing in terms of the RMSE for both the logarithmically transformed implied volatility (the first two columns) and the level of implied volatility (the last two columns). Interestingly, the Auxiliary structure only results in slightly larger pricing errors than the DHNG model. This can be ascribed to the adaptive nature of the score-driven model, which shifts $\log\eta_{t}$ to a higher level to compensate for errors embedded in the Auxiliary pricing formulae. Thus, while the first-order approximation is insufficient, the estimated score-drive model “rescues” the Auxiliary structure, by inflating the variance risk ratio above its true value.
Table S.3: Estimation Results for Auxiliary and DHNG
@@
Table S.4: Pricing Performance for Auxiliary and DHNG
@@
S.4 Supplementary Material for Table 6
@@
Economic Policy Uncertainty 0.380*** 0.208***
@@
Survey-based Uncertainty 0.148*** 0.078***
@@
Economic Uncertainty Index 0.071*** 0.013
@@
Variance Risk Premium 0.120*** 0.096***
@@
Stock Market Volatility 0.394*** 0.395*** 0.368*** 0.356*** 0.294*** 0.363*** 0.354*** 0.395*** 0.306***
@@
Default Yield Spread 0.385*** 0.358*** 0.372*** 0.365*** 0.393*** 0.282*** 0.246*** 0.285*** 0.196***
@@
Default Return Spread 0.026*** 0.027*** 0.025*** 0.025*** 0.019*** 0.024*** 0.021*** 0.026*** 0.020***
@@
Long Term Yield 0.028** 0.030** 0.040*** 0.043*** 0.077*** 0.016 0.045*** 0.035*** 0.063***
@@
Dividend Price Ratio -0.457*** -0.435*** -0.525*** -0.529*** -0.638*** -0.463*** -0.403*** -0.354*** -0.471***
@@
Earnings Price Ratio -0.080 -0.079 -0.089 -0.056 -0.096 -0.094 0.086 -0.015 -0.007
@@
Book to Market Ratio 0.484 0.149 0.574 0.485 0.402 0.625 0.023 0.230 -0.095
@@
Long Term Rate of Return 0.002 0.003 0.004 0.004 0.001 0.003 0.004 0.006 0.006*